# Dialect Matters: Cross-Lingual ASR Transfer for Low-Resource Indic Language Varieties

### v1

@@ 1 Introduction
    We introduce a general framework for quantifying bias toward pre-training languages in dialectal ASR, enabling systematic analysis of model behavior on non-standardized and code-mixed speech.
@@ 2 Related Work
Recent work has investigated improving speech recognition for Indic dialects and language varieties. For instance, Kumar et al. (2025) evaluate a multilingual dialect identification and ASR system that combines Conformer-based speech representations with a RoBERTa model across 33 dialects of 8 Indic languages using read speech. Related efforts have also aimed to develop comprehensive ASR toolkits for Indic languages Singh Chadha et al. (2022). A similar study on standard to dialectal transfer learning in German language has been done Blaschke et al. (2025). However, cross-dialectal ASR adaptation across Indic languages, particularly in zero-shot settings and on spontaneous speech, has not yet been systematically studied, which we address in this work.
@@
The widespread presence of multiple language varieties in India, shaped by historical and sociolinguistic factors, has led to extensive code-mixing in everyday speech, posing additional challenges for ASR systems. While code-mixed speech has been studied in specific settings such as Hindi-Marathi ASR Palivela et al. (2025), we propose a framework for quantifying transcription errors arising from code-mixing in dialectal ASR.
@@ 3 Languages and Data
The Indic linguistic landscape is characterized by a high degree of diversity, encompassing hundreds of languages and dialects with varying levels of standardization and resource availability. Many widely spoken languages coexist with numerous regional dialects that differ substantially in phonology, morphology, and lexicon, despite phylogenetic relatedness. In everyday use, speakers frequently engage in code-mixing, particularly with English, and in spontaneous and acoustically noisy environments. These factors pose major challenges for ASR systems trained primarily on clean, standardized, and monolingual speech. We present a phylogenetic tree of the languages included in our experiments in Figure 1.
@@
VAANI consists of spontaneous speech collected by prompting participants to describe images in their local dialect. The dataset contains over 150,000 hours of speech, of which approximately 10% is transcribed, and covers speakers from 773 districts across India. As Indic languages are written in multiple scripts, we focus our analysis on language varieties written in the Devanagari script. Due to the broad coverage of dialects and language varieties, the amount of available data varies substantially across languages in the dataset; we account for this variability as much as possible in our experimental design.
@@
Beyond conducting cross-lingual experiments across Indic language varieties, we also conduct a detailed analysis of ASR for one nonstandard language variety. We select Garhwali for this case study. [^2] According to the Post-1971 Census of India, these languages were recognized as “mother tongues” and were designated as dialects of Hindi Khubchandani (1991); however, some linguists argue that Garhwali and other Pahari languages are distinct languages Gusain et al. (2023). Garhwali is in the Pahari subgroup [^3] We refer to the lowest non-leaf nodes in a phylogenetic tree as subgroups. spoken in the Himalayan region, particularly in the state of Uttarakhand Gusain et al. (2023).
@@ 4 Experiments and Results
In our experiments, we aim to answer the following questions:
@@
4.1 Analyzing Broader Patterns in Indic ASR
RQ1: Baseline Assessment
@@ Results
RQ3: Cross-Lingual Transfer
We wanted to examine whether there are any performance gains from cross-dialectal fine-tuning in zero-shot settings. For each of these experiments, we fine-tune w2v-bert-2.0555https://huggingface.co/facebook/w2v-bert-2.0 (selected through preliminary comparison of models) on 1 to 7 hours of speech, depending on the availability. Each model was then evaluated on the test set of all Devanagari script languages available from VAANI. We compute correlations between the error rate and phylogenetic distance under two conditions: across all evaluation languages and restricted to nonstandard test varieties. We additionally measure the association between the error rate and whether the fine-tuning language represents dialectal or standard speech. We consider the following languages from the dataset to be dialects or nonstandard varieties: Angika Awadhi, Bajjika, Bhili, Chhattisgarhi, Garhwali, Halbi, Haryanvi, Jaipuri, Khariboli, Khortha, Kumaoni, Magadhi, Malvani, Marwari, Nagpuri, Surjapuri, and Thethi. [^6] None of these languages belong to the Eighth Schedule to the Indian Constitution listing the officially recognized languages (https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India).
@@
Dialectal Transfer Effects
@@
Fine-Tuning Lang Hours Jaipuri
@@
Marwari only (dialect) 17.2 0.301 0.162
@@
Fine-Tuning Lang Hours Thethi
@@
Fine-Tuning Lang Hours Haryvani
@@
4.2 Garhwali ASR Case Study
Baseline Selection
@@ 5 Conclusion
We further present the first detailed ASR analysis for Garhwali, a nonstandard Pahari language variety, and show that a w2vBERT-based model fine-tuned on Garhwali achieves the best performance among the evaluated architectures. Although the resulting word error rate of 49.3% remains insufficient for fully automated transcription, this case study illustrates both the challenges of dialectal ASR and the benefits of dialect-specific modeling.
Finally, we introduce a quantitative framework for analyzing bias toward pre-training languages in dialectal ASR. Our error analysis reveals substantial bias toward Hindi in both multilingual and Hindi-fine-tuned models, manifesting in systematic normalization of dialectal and code-mixed forms. This framework underscores the need for dialect-aware data selection and modeling strategies in future ASR systems.

### v2

@@ 1 Introduction
    We introduce a diagnostic approach for quantifying bias toward pre-training languages in dialectal ASR, enabling systematic analysis of model behavior on non-standardized and code-mixed speech.
@@ 2 Related Work
Recent work has highlighted systematic performance gaps between standard language varieties and regional or minority dialects across NLP and speech technologies. Kantharuban et al. (2023) provide a large-scale evaluation of state-of-the-art models for machine translation and automatic speech recognition across regional dialects of multiple high- and low-resource languages, showing that dialectal performance disparities are widespread and variably correlated with linguistic, social, and data-related factors. Complementary to this line of work, Blaschke et al. (2025) study standard-to-dialect transfer in spoken and written settings for German, demonstrating that speech-based models are more robust to dialectal variation than text-based or cascaded approaches, particularly when orthographic normalization is involved.
In the Indic context, prior work has investigated dialectal variation primarily through multilingual or language-specific ASR systems. For example, Kumar et al. (2025) evaluate a multilingual dialect identification and ASR pipeline across 33 dialects of eight Indic languages using read speech, while other efforts have focused on building comprehensive ASR toolkits for standardized Indic languages Singh Chadha et al. (2022). However, cross-dialectal and cross-lingual transfer for Indic ASR—especially in zero-shot settings and on spontaneous, non-standardized speech—has not yet been systematically examined. Our work addresses this gap by analyzing cross-lingual transfer behavior and dialectal bias in low-resource Indic language varieties.
@@
The widespread presence of multiple language varieties in India, shaped by historical and sociolinguistic factors, has led to extensive code-mixing in everyday speech, posing additional challenges for ASR systems. While code-mixed speech has been studied in specific settings such as Hindi-Marathi ASR Palivela et al. (2025), we provide a systematic way to quantify transcription errors arising from code-mixing in dialectal ASR.
@@ 3 Languages and Data
The Indic linguistic landscape is characterized by a high degree of diversity, encompassing hundreds of languages and dialects with varying levels of standardization and resource availability. Many widely spoken languages coexist with numerous regional dialects that differ substantially in phonology, morphology, and lexicon, despite phylogenetic relatedness Massica (1993). In everyday use, speakers frequently engage in code-mixing, particularly with English, and in spontaneous and acoustically noisy environments. These factors pose major challenges for ASR systems trained primarily on clean, standardized, and monolingual speech Diwan et al. (2021). We present a phylogenetic tree of the languages included in our experiments in Figure 1.
@@
VAANI consists of spontaneous speech collected by prompting participants to describe images in their local dialect. The dataset contains over 150,000 hours of speech, of which approximately 10% is transcribed, and covers 156,534 speakers from 773 districts across India. As Indic languages are written in multiple scripts, we focus our analysis on language varieties written in the Devanagari script. Due to the broad coverage of dialects and language varieties, the amount of available data varies substantially across languages in the dataset; we account for this variability as much as possible in our experimental design.
@@
Beyond conducting cross-lingual experiments across Indic language varieties, we also conduct a detailed analysis of ASR for one nonstandard language variety. We select Garhwali for this case study. [^2] According to the Post-1971 Census of India, these languages were recognized as “mother tongues” and were designated as dialects of Hindi Khubchandani (1991); however, some linguists argue that Garhwali and other Pahari languages are distinct languages Gusain et al. (2023). Garhwali is a language that belongs to the Pahari subgroup  [^3] We refer to the lowest non-leaf nodes in a phylogenetic tree as subgroups. spoken in the Himalayan region, particularly in the state of Uttarakhand Gusain et al. (2023).
@@ 4 Experiments and Results
In our experiments, we aim to answer the following questions to analyze broader patterns in Indic ASR:
@@
4.1 RQ1: Baseline Assessment for Indic ASR
Setup
@@ Results
4.3 RQ3: Cross-Lingual Transfer
We wanted to examine whether there are any performance gains from cross-dialectal fine-tuning in zero-shot settings. For each of these experiments, we fine-tune w2v-bert-2.0555https://huggingface.co/facebook/w2v-bert-2.0 (selected through preliminary comparison of models) on 1 to 7 hours of speech, depending on the availability. Each model was then evaluated on the test set of all Devanagari script languages available from VAANI. We compute correlations between the error rate and phylogenetic distance under two conditions: across all evaluation languages and restricted to nonstandard test varieties. We additionally measure the association between the error rate and whether the fine-tuning language represents dialectal or standard speech. We consider the following languages from the dataset to be dialects or nonstandard varieties: Angika, Awadhi, Bajjika, Bhili, Chhattisgarhi, Garhwali, Halbi, Haryanvi, Jaipuri, Khariboli, Khortha, Kumaoni, Magadhi, Malvani, Marwari, Nagpuri, Surjapuri, and Thethi. [^6] None of these languages belong to the Eighth Schedule to the Indian Constitution listing the officially recognized languages (https://en.wikipedia.org/wiki/Eighth_Schedule_to_the_Constitution_of_India).
@@
4.3.1 Dialectal Transfer Effects
@@
Fine-Tuning Language Hours Jaipuri
@@
Marwari (dialect only) 17.2 0.301 0.162
@@
Fine-Tuning Language Hours Thethi
@@
Fine-Tuning Language Hours Haryvani
@@
5 Garhwali ASR Case Study
5.1 Baseline Selection
Setup
@@ 6 Conclusion
We further present the first detailed ASR analysis for Garhwali, a nonstandard Pahari language variety, and show that a w2vBERT-based model fine-tuned on Garhwali achieves the best performance among the evaluated architectures. Although the resulting word error rate of 49.3% remains insufficient for fully automated transcription, this case study illustrates both the challenges of dialectal ASR and the benefits of dialect-specific modeling. Our quantitative error analysis further reveals substantial bias toward Hindi in both multilingual and Hindi-fine-tuned models, manifesting in systematic normalization of dialectal and code-mixed forms, and underscoring the need for dialect-aware data selection and modeling strategies in future ASR systems.